OpenStudy (anonymous):

Find a function y = f(x) whose second derivative is y" =12x - 2 at each point (x, y) on its graph and y= -x + 5 is tangent to the graph at the point corresponding to x = 1 .

OpenStudy (cwrw238):

integrate 12x - 2 to gety' y'= 6x^2 - 2x + c integrate again to get y y = 2x^3 - x^2 + cx + d where c and d are constants

OpenStudy (cwrw238):

the slope of the tangent = f' = 6x^2 - 2x + c at x = 1 this is 6 - 2 + c the tangent equation is y = -x + 5 so slope = -1 = 6-2+c giving c = -1 -6+ 2 = -5 giving f' = 6x^2 -2x -5 and fx) = 2x^3 -x^2 - 5x + d now we need to find value of d

OpenStudy (cwrw238):

tangent equation is y = -x + 5 when x = 1 y = -1 + 5 = 4 this point is also on graph of f(x) so 4 = 2(1)^3 -1^2 - 5(1) + d solve this for d and plug this into the expression we got in last post